OpenStudy (anonymous):

if (x)=4x, g(x)=5x-3, and h(x)=√x, find (f o g o h)(4)

OpenStudy (anonymous):

fog[√4]=fog[2]=f[5*2-3]=f[7]=4*7=28

OpenStudy (anonymous):

so 28 is the answer

OpenStudy (anonymous):

Yeap